# On the Definition of Intelligence

### v1

@@ Abstract
To engineer AGI, we should first capture the essence of intelligence in a species-agnostic form that can be evaluated, while being sufficiently general to encompass diverse paradigms of intelligent behavior, including reinforcement learning, generative models, classification, analogical reasoning, and goal-directed decision-making. We propose a general criterion based on sample fidelity: intelligence is the ability, given sample(s) from a category, to generate sample(s) from the same category. We formalise this intuition as $\varepsilon$-category intelligence: it is $\varepsilon$-intelligent with respect to a category if no chosen admissible distinguisher can separate generated from original samples beyond tolerance $\varepsilon$. We present the formal framework, outline empirical protocols, and discuss implications for evaluation, safety, and generalization. By defining intelligence based on the principle of generative fidelity to a category, our definition provides a single yardstick for comparing biological, artificial, and hybrid systems, and invites further theoretical refinement and empirical validation.
@@ 1.1 Motivation
We boldly propose to adopt a certain property of analogy as the foundational definition of intelligence. Prior research has emphasized the centrality of analogy in cognition[13]. We present this property as an informal definition in Section 2, and demonstrate how key aspects such as creativity and learning ability can be situated within this framework. To formally incorporate relational and structural aspects into the definition, we adopt a precise mathematical notation to capture the similarity between samples. In this setting, samples within a category exhibit a form of similarity, often of an abstract nature. To evaluate such similarity, we define $\varepsilon$-category intelligence using precise mathematical terminology, allowing us to concretely measure intelligence through the notion of indistinguishability, rather than relying on vague descriptions. Ultimately, we provide a new lens that stimulates further discussion and refinement on the nature of intelligence.
@@ 2 Proposed Definition
Intelligence is the ability, given sample(s) from a category, to generate sample(s) from the same category.
Let $X$ denote the universe of all conceivable samples. A category space is a surjective mapping
$\Phi:X\longrightarrow K,$
where $K$ is the set (or measurable space) of category labels. For any label $k\in K$ we write
$C_{k}\;=\;\Phi^{-1}(k)\;\subseteq\;X,$
and call $C_{k}$ the category fibre associated with $k$.
The given samples form a finite multiset [^1] Duplicates are permitted. Whenever order is irrelevant we view $S$ as a finite multiset $S\in M(X)$; when order is relevant (e.g. for sequence-based distinguishers) we instead treat $S=(x_{1},\dots,x_{m})\in X^{m}$ as a finite sequence.
$S\;=\;\{x_{1},\dots,x_{m}\}\subseteq X,$
@@
Definition 1 ($\varepsilon$-Category Intelligence)
[…]
@@
The ability to generate set $\widehat{S}$ is said to be $\varepsilon$-intelligent with respect to the (implicit) category $C$ if $\Delta_{F}^{\sigma}(S,\widehat{S})\leq\varepsilon$.
Informally, $\widehat{S}$ is indistinguishable from the given samples $S$—up to tolerance $\varepsilon$—for every distinguisher in $F$, according to the chosen scoring rule $\sigma$. The principled selection of categories $C$ and distinguishers $F$ is a non-trivial question on which we will elaborate in Section 5.
For example, consider ChatGPT producing Studio Ghibli-style frames: let $S$ be a set of genuine Ghibli images and $\widehat{S}$ be ChatGPT’s outputs. If human judges (or automated critics) cannot reliably distinguish $\widehat{S}$ from $S$ within a small tolerance $\varepsilon$, then by our definition ChatGPT is $\varepsilon$-intelligent in the category “Ghibli-style images.” This scenario aligns with our intuitive notion of “looks like Ghibli,” illustrating how indistinguishability underpins both formal measurement and everyday understanding of style-based intelligence.
Crucially, the similarity among samples in a category can be highly abstract. For instance, a horse and a car, despite their stark differences in appearance and internal structure, belong to the same category of “conveyances for human transport.” An entity demonstrates intelligence with respect to this category if, given the concept of a horse, it can generate the concept of a car or other novel solutions for transport.
@@
  • Operational: Intelligence can be measured through categorical alignment between observable samples and generated samples.
  • Falsifiable: Systems that consistently generate outputs divergent from the reference category exhibit low intelligence.
@@
Samples are not limited to data points—they can also include products, concepts, designs, actions, and more. Identifying a category $C$ whose distinguishers $F$ cannot separate from the data at hand is itself an act of intelligence.
@@ 3 Interpretations and Special Cases
3.1 The Turing Test as a Special Case of $\varepsilon$-Category Intelligence
If we define the category $C$ as the set of human language responses and let the family of distinguishers $\mathcal{F}$ consist of human judges evaluating whether a given response was produced by a human or a machine, then the classic Turing Test naturally arises as a special case of $\varepsilon$-category intelligence, with respect to electronic computers and humans.
3.2 C-test as a Special Case of $\varepsilon$-Category Intelligence
Following Hernández-Orallo’s C-tests [12], we define, for each difficulty level $h$, the category $C_{h}$ as the set of all sequences whose Levin complexity equals a fixed value $h$ and that admit a single minimal explanation. Let the distinguisher family $F$ accept only the continuation produced by this unique minimal program. Under this configuration, the $\varepsilon$-criterion degenerates into a binary verdict: if an agent’s continuation matches the unique minimal explanation, then $\Delta_{F}^{\sigma}(S,S^{\prime})\leq\varepsilon$ and the response is correct; otherwise it is incorrect. Hence the classical C-test can be regarded as an $\varepsilon$-Category instance with a vanishingly small $\varepsilon$.
3.3 Legg–Hutter Intelligence as a Facet of $\varepsilon$-Category Intelligence
If we define each category $C$ as the set of behaviour sequences that maximise expected reward in a particular environment and let the family of distinguishers $F$ be the environment’s reward signal judging how closely a generated behaviour approaches that optimum, then, when we consider only an agent interacting with its environment, Legg–Hutter intelligence naturally arises as a facet of $\varepsilon$-category intelligence.
@@ 4 Foundational Properties of the Definition
4.1 Incompleteness and Transformation of Samples
The given samples may be incomplete or transformed. All forms of intelligence can be seen as methods for generating new samples based on given samples. For example, a classic neural network trained on a dataset learns to approximate a surface that locally aligns with the target category, thereby enabling consistent sample generation. Human intelligence is another method. There will be future machine learning methods that do not rely solely on surface approximation to generate samples.
4.2 Categorical Ambiguity of Samples
A single set of samples may correspond to multiple possible categories. An observable sample can often belong to multiple categories. The initial sample of a category is often generated from another category, which can be understood as a form of creativity. Understanding how samples relate to their generative processes—and transitions between categories—is a profound and ongoing research problem.
@@ 4.3 The Locality of Intelligence
[…]
@@ 4.4 Generation as Core Expression of Intelligence
Our definition emphasizes that intelligence is not merely a matter of abstract mapping or numeric manipulation—as is often the case in formal mathematics—but fundamentally about generation. The act of generating new samples from a given category lies at the heart of intelligent behavior.
@@
Viewed through this lens, intelligence, defined as the ability to generate category-consistent samples, appears to be a principle widely present in natural systems. The field of bio-inspired engineering, where designs mimic biological structures and processes (e.g., shark skin-inspired surfaces for drag reduction, or ant colony optimization for algorithms), can be seen as implicitly leveraging this principle [2, 1, 6]. These biological “designs” or behaviors represent highly effective solutions within specific environmental or functional categories. Thus, bio-mimicry can be interpreted as generating samples (structures, behaviors) that are highly consistent with the success criteria of their respective ecological categories.
The architectural wisdom of I.M. Pei provides an elegant analogy. He once remarked: “There is always a theme, there is a certain repetition, but they do not seem like repetition—only the endless variety of a simple theme.”[19] This resonates deeply with our formulation, in which intelligence expresses itself through variety constrained by coherence—a generative consistency within categorical identity.
@@ 4.5 Intelligence and Consciousness
As a counterexample, consider a minimal computer program designed to solve calculus problems through symbolic manipulation or numerical methods. Within the category $C=\{\text{correctly solved calculus problems}\}$, this program can exhibit extremely high $\varepsilon$-category intelligence, producing outputs (i.e., solutions) that are indistinguishable from valid mathematical results—likely surpassing the performance of most humans in this narrow domain. Yet few would attribute consciousness to such a deterministic mathematical solver, as its algorithmic nature is generally not regarded as a sufficient condition for consciousness.
@@ 4.6 Intelligence and Learning Ability
It is frequently assumed that intelligence necessitates an inherent capacity for learning. Our proposed definition, however, does not mandate this. We can illustrate this distinction with a counterexample within our framework. Suppose a standard LLM is capable of consistently generating correct answers ($\widehat{S}$) for problems belonging to a specific category $C$. Now, we fix its internal parameters and restrict its computational process to be purely deterministic. Such a system is clearly incapable of learning, as it cannot update its internal state based on new information or interactions, but according to our definition, it is intelligent within that domain. Clearly, learning ability is related to the change in $\Delta_{F}^{\sigma}(S,\widehat{S})$ instead.
@@ 4.7 Clarifications on Scope and Interpretation
  • Classification: In classification tasks, each data point represents an input-output pair from a joint category. A system that accurately produces such pairs aligns with the category and is thus intelligent by this definition.
  • Decision-Making: In decision-making or interactive environments, the “category” refers to the set of actions likely to achieve a given objective (e.g., survival, success, reward maximization). Producing a sample corresponds to selecting a behavior consistent with the category of goal-achieving decisions.
  • Memory-Based Methods: Systems that rely heavily on memorization or exhaustive enumeration can still qualify as intelligent if their outputs remain consistent with the target category, although they may lack generality or efficiency.
  • Thought Experiment: Consider a person who has memorized all knowledge of Wikipedia and answers questions accordingly, compared to someone who cannot. In the category of Wikipedia-based question answering, the former is demonstrably more intelligent. The significance of memorizing samples from a category is often underestimated. If a person were capable of perfectly memorizing the answers to all questions within a domain, they would, by definition, exhibit a high degree of intelligence—at least within that domain.
@@ 5 A Dynamic Framework for General Intelligence
5.1 On the Selection of Categories and Distinguishers
The operationalization of our framework hinges on the principled selection of a category $C$ and its corresponding family of distinguishers $F$. These choices are governed by pragmatic fitness and measurement resolution, rather than being arbitrary.
First, the selection of a category $C$ is not a normative judgement of its intrinsic “rightness” or “goodness.” Instead, this constitutes a process of pragmatic selection: categories that effectively satisfy human needs become more prevalent because of their fitness, and are thus more frequently observed and studied. More broadly, human needs represent merely one possible selective environment. In a more general sense, the prominence of any category arises from its fitness within a given environment, be it physical, ecological, or an abstract logical system.
Second, given a category $C$ and a target intelligence level $\varepsilon$, the choice of the distinguisher family $F$ is constrained. Any practical $F$ has its own intrinsic resolution limit, denoted as $\varepsilon_{f}$—its margin of error in determining membership in $C$. A fundamental prerequisite for a meaningful assessment is that $\varepsilon_{f}<\varepsilon$, as an instrument cannot resolve details finer than its own precision. This imposes a critical limitation: if an agent’s performance ($\varepsilon$) far exceeds the evaluator’s capability, such that $\varepsilon$ is significantly smaller than $\varepsilon_{f}$, then $F$ can no longer reliably verify such high-level performance.
@@ 5.2 Efficiency, Cost, and Dynamic Adaptation
In principle, within a system that evolves over time, an ideal general intelligent entity should possess the potential to generate samples for any category that can be specified or exemplified. In reality, however, any entity at any single point in time $t$ confronts a finite set of categories, $C_{t}$. It is this very transition from infinite potential to finite reality that necessitates evaluating intelligence from a dynamic perspective. From this vantage point, intelligence can be assessed along two dimensions: [^2] The terms “synchronic” and “diachronic” are adopted from structural linguistics to distinguish the study of a phenomenon at a particular point in time from its evolution through time, respectively. See Saussure (1916) [22].
@@
[…]
@@
    Diachronic Capability: The ability of an entity to transition from a state of competence for a set of categories $C_{t}$ to a subsequent set $C_{t^{\prime}}$. This essentially measures the entity’s adaptive capacity.
Within this framework, concepts such as efficiency, cost, compression, and learning acquire their instrumental value, their importance realized through their potential contribution to these two capabilities. Whether and to what extent an entity leverages these properties stems from the nature of the challenges it confronts—namely, the scale, complexity, and temporal dynamics of the category sets $C_{t}$.
For instance, when confronted with a large and diverse $C_{t}$, a low-cost generation process might be a strategy for enhancing synchronic capability. Likewise, when category sets change frequently and unpredictably, a highly compressed knowledge representation could become an advantage for boosting diachronic capability. These are not universal laws, however; the ultimate determinant of a strategy’s merit remains the entity’s overall ability to sustainably generate category samples.
6 Generalization to Unseen Categories
We now consider a training setup based on multiple categories. Let the training data consist of samples:

### v2

@@ Abstract
To engineer AGI, we should first capture the essence of intelligence in a species-agnostic form that can be evaluated, while being sufficiently general to encompass diverse paradigms of intelligent behavior, including reinforcement learning, generative models, classification, analogical reasoning, and goal-directed decision-making. We propose a general criterion based on entity fidelity: Intelligence is the ability, given entities exemplifying a concept, to generate entities exemplifying the same concept. We formalise this intuition as $\varepsilon$-concept intelligence: it is $\varepsilon$-intelligent with respect to a concept if no chosen admissible distinguisher can separate generated entities from original entities beyond tolerance $\varepsilon$. We present the formal framework, outline empirical protocols, and discuss implications for evaluation, safety, and generalization. By defining intelligence based on the principle of generative fidelity to a concept, our definition provides a single yardstick for comparing biological, artificial, and hybrid systems, and invites further theoretical refinement and empirical validation.
@@ 1.1 Motivation
We boldly propose to adopt a certain property of analogy as the foundational definition of intelligence. Prior research has emphasized the centrality of analogy in cognition[13]. We present this property as an informal definition in Section 2, and demonstrate how key aspects such as creativity and learning ability can be situated within this framework. To formally incorporate relational and structural aspects into the definition, we adopt a precise mathematical notation to capture the similarity between entities. In this setting, entities within a concept exhibit a form of similarity, often of an abstract nature. To evaluate such similarity, we define $\varepsilon$-concept intelligence using precise mathematical terminology, allowing us to concretely measure intelligence through the notion of indistinguishability, rather than relying on vague descriptions. Ultimately, we provide a new lens that stimulates further discussion and refinement on the nature of intelligence.
@@ 2 Proposed Definition
Intelligence is the ability, given entities exemplifying a concept, to generate entities exemplifying the same concept.
Let $U$ denote the universe of all conceivable entities. Entities are not limited to data points—they can also include products, designs, actions, or even concepts and more. A conceptual space is a surjective mapping
$\Phi:U\longrightarrow K,$
where $K$ is the set (or measurable space) of concepts, each $k\in K$ serving as a representation for a specific set of entities in $U$. For any concept $k\in K$ we write
$C_{k}\;=\;\Phi^{-1}(k)\;\subseteq\;U,$
and call $C_{k}$ the fibre associated with $k$.
The given entities form a finite multiset [^1] Duplicates are permitted. Whenever order is irrelevant, we view $S$ as a finite multiset $S\in M(U)$; when order is relevant (e.g., for sequence-based distinguishers), we instead treat $S=(x_{1},\dots,x_{m})\in U^{m}\text{ as a finite sequence.}$
$S\;=\;\{x_{1},\dots,x_{m}\}\subseteq U,$
@@
Definition 1 ($\varepsilon$-Concept Intelligence)
Let $F$ be a family of distinguishers $f:U\!\to\![0,1]$. Let $S\subseteq U$ be a given entity multiset that is compatible with a concept (i.e., there exists at least one concept $k\in K$ such that its fibre $C_{k}$ contains—or is well-approximated by—$S$). Given any generated entity set $\widehat{S}\subseteq U$ produced on the basis of $S$, and scoring function $\sigma:M([0,1])\longrightarrow[0,1]$, define
@@
The ability to generate set $\widehat{S}$ is said to be $\varepsilon$-intelligent with respect to the (implicit) fibre $C_{k}$ if $\Delta_{F}^{\sigma}(S,\widehat{S})\leq\varepsilon$.
Informally, $\widehat{S}$ is indistinguishable from the given entities SS—up to tolerance $\varepsilon$—for every distinguisher in $F$, according to the chosen scoring rule $\sigma$. The principled selection of a concept $k$ and distinguishers $F$ is a non-trivial question on which we will elaborate in Section 5.
For example, consider ChatGPT producing Studio Ghibli-style frames: let $S$ be a set of genuine Ghibli-style images and $\widehat{S}$ be ChatGPT’s outputs. If human judges (or automated critics) cannot reliably distinguish $\widehat{S}$ from $S$ within a small tolerance $\varepsilon$, then by our definition ChatGPT is $\varepsilon$-intelligent in the concept “Ghibli-style images.” This scenario aligns with our intuitive notion of “looks like Ghibli,” illustrating how indistinguishability underpins both formal measurement and everyday understanding of style-based intelligence.
Crucially, the similarity among entities in a concept can be highly abstract. For instance, a horse and a car, despite their stark differences in appearance and internal structure, belong to the same concept of “conveyances for human transport.” An entity demonstrates intelligence with respect to this concept if, given an entity (e.g., a horse) exemplifying it, it can generate another entity (e.g., a car) that also exemplifies it, or other novel solutions for transport.
@@
  • Operational: Intelligence can be measured through conceptual alignment between observable entities and generated entities.
  • Falsifiable: Systems that consistently generate outputs divergent from the reference concept exhibit low intelligence.
@@
Identifying a concept $k$ whose distinguishers $F$ cannot separate from the data at hand is itself an act of intelligence.
@@ 3 Interpretations and Special Cases
3.1 The Turing Test as a Special Case of $\varepsilon$-Concept Intelligence
If we define the fibre $C_{k}$ (corresponding to a concept $k$) as the set of human language responses and let the family of distinguishers $F$ consist of human judges evaluating whether a given response was produced by a human or a machine, then the classic Turing Test naturally arises as a special case of $\varepsilon$-concept intelligence, with respect to electronic computers and humans.
3.2 C-test as a Special Case of $\varepsilon$-Concept Intelligence
Following Hernández-Orallo’s C-tests [12], we define, for each difficulty level $h$, the fibre $C_{k_{h}}$ (corresponding to a concept $k_{h}$) as the set of all sequences whose Levin complexity equals a fixed value $h$ and that admit a single minimal explanation. Let the distinguisher family $F$ accept only the continuation produced by this unique minimal program. Under this configuration, the $\varepsilon$-criterion degenerates into a binary verdict: if an agent’s continuation matches the unique minimal explanation, then $\Delta_{F}^{\sigma}(S,\widehat{S})\leq\varepsilon$ and the response is correct; otherwise it is incorrect. Hence the classical C-test can be regarded as an $\varepsilon$-Concept instance with a vanishingly small $\varepsilon$.
3.3 Legg–Hutter Intelligence as a Facet of $\varepsilon$-Concept Intelligence
If we define the fibre $C_{k}$ (corresponding to a concept $k$) as the set of behaviour sequences that maximise expected reward in a particular environment and let the family of distinguishers $F$ be the environment’s reward signal judging how closely a generated behaviour approaches that optimum, then, when we consider only an agent interacting with its environment, Legg–Hutter intelligence naturally arises as a facet of $\varepsilon$-concept intelligence.
@@ 4 Foundational Properties of the Definition
4.1 Incompleteness and Transformation of Entities
The given entities may be incomplete or transformed. All forms of intelligence can be seen as methods for generating new entities based on given entities. For example, a classic neural network trained on a dataset learns to approximate a surface that locally aligns with the target concept, thereby enabling consistent entity generation. Human intelligence is another method. There will be future machine learning methods that do not rely solely on surface approximation to generate entities.
4.2 Conceptual Ambiguity of Entities
A single set of entities may correspond to multiple possible concepts. An observable entity can often belong to multiple concepts. The initial entity exemplifying a concept is often generated as an instance of another concept, which can be understood as a form of creativity. Understanding how entities relate to their generative processes—and transitions between concepts—is a profound and ongoing research problem.
@@ 4.3 The Locality of Intelligence
Intelligence, as defined here, is local to the given concept. For example, while a homing pigeon cannot code or speak human language, after just 15–30 days of training it can classify breast-cancer histopathology slides with up to $95\%$ accuracy, and an ensemble of four pigeons achieves an AUC of 0.990.99—on par with specialist pathologists[17]. Likewise, if a machine consistently translates languages more accurately than a person, it is more intelligent than a person in language translation. More general intelligence would imply the ability to generate entities across a wider range of concepts.
@@ 4.4 Generation as Core Expression of Intelligence
Our definition emphasizes that intelligence is not merely a matter of abstract mapping or numeric manipulation—as is often the case in formal mathematics—but fundamentally about generation. The act of generating new entities exemplifying a given concept lies at the heart of intelligent behavior.
@@
Viewed through this lens, intelligence, defined as the ability to generate concept-consistent entities, appears to be a principle widely present in natural systems. The field of bio-inspired engineering, where designs mimic biological structures and processes (e.g., shark skin-inspired surfaces for drag reduction, or ant colony optimization for algorithms), can be seen as implicitly leveraging this principle [2, 1, 6]. These biological “designs” or behaviors represent highly effective solutions within specific environmental or functional concepts. Thus, bio-mimicry can be interpreted as generating entities (structures, behaviors) that are highly consistent with the success criteria of their respective ecological concepts.
The architectural wisdom of I.M. Pei provides an elegant analogy. He once remarked: “There is always a theme, there is a certain repetition, but they do not seem like repetition—only the endless variety of a simple theme.”[19] This resonates deeply with our formulation, in which intelligence expresses itself through variety constrained by coherence—a generative consistency within conceptual identity.
@@ 4.5 Intelligence and Consciousness
As a counterexample, consider a minimal computer program designed to solve calculus problems through symbolic manipulation or numerical methods. Within the conceptual framework where the fibre $C_{k}$ (corresponding to a concept $k$) is the set of correctly solved calculus problems, this program can exhibit extremely high $\varepsilon$-concept intelligence, producing outputs (i.e., solutions) that are indistinguishable from valid mathematical results—likely surpassing the performance of most humans in this narrow domain. Yet few would attribute consciousness to such a deterministic mathematical solver, as its algorithmic nature is generally not regarded as a sufficient condition for consciousness.
@@ 4.6 Intelligence and Learning Ability
It is frequently assumed that intelligence necessitates an inherent capacity for learning. Our proposed definition, however, does not mandate this. We can illustrate this distinction with a counterexample within our framework. Suppose a standard LLM is capable of consistently generating correct answers ($\widehat{S}$) for problems belonging to a specific concept $k$. Now, we fix its internal parameters and restrict its computational process to be purely deterministic. Such a system is clearly incapable of learning, as it cannot update its internal state based on new information or interactions, but according to our definition, it is intelligent within that domain. Clearly, learning ability is related to the change $\Delta_{F}^{\sigma}(S,\widehat{S}) in$stead.
@@ 4.7 Clarifications on Scope and Interpretation
  • Classification: In classification tasks, each input-output pair is an entity. A system that accurately produces such entities, exemplifying a joint concept, aligns with that concept and is thus intelligent by this definition.
  • Decision-Making: In decision-making or interactive environments, we identify a concept (e.g., “goal-achieving behaviors”) whose fibre is the set of behaviors (which are entities) likely to achieve a given objective (e.g., survival, success, reward maximization). Producing an entity here corresponds to selecting a behavior consistent with this concept.
  • Memory-Based Methods: Systems that rely heavily on memorization or exhaustive enumeration can still qualify as intelligent if their outputs remain consistent with the target concept. The significance of memorizing entities from a concept is often underestimated.
@@ 5 A Dynamic Framework for General Intelligence
5.1 On the Selection of Concepts and Distinguishers
The operationalization of our framework hinges on the principled selection of a concept $k$ (and its corresponding fibre $C_{k}$) and its family of distinguishers $F$. These choices are governed by pragmatic fitness and measurement resolution, rather than being arbitrary.
First, the selection of a concept $k$ (and its corresponding fibre $C_{k}$) is not a normative judgement of its intrinsic “rightness” or “goodness.” Instead, this constitutes a process of pragmatic selection: concepts whose fibres effectively satisfy human needs become more prevalent because of their fitness, and are thus more frequently observed and studied. More broadly, human needs represent merely one possible selective environment. In a more general sense, the prominence of any concept (and its fibre) arises from its fitness within a given environment, be it physical, ecological, or an abstract logical system.
Second, given a concept $k$ (and its fibre $C_{k}$) and a target intelligence level $\varepsilon$, the choice of the distinguisher family $F$ is constrained. Any practical $F$ has its own intrinsic resolution limit, denoted as $\varepsilon_{f}$—its margin of error in determining membership in $C_{k}$. A fundamental prerequisite for a meaningful assessment is that $\varepsilon_{f}<\varepsilon$, as an instrument cannot resolve details finer than its own precision. This imposes a critical limitation: if an agent’s performance ($\varepsilon$) far exceeds the evaluator’s capability, such that $\varepsilon$ is significantly smaller than $\varepsilon_{f}$, then $F$ can no longer reliably verify such high-level performance.
@@ 5.2 Efficiency, Cost, and Dynamic Adaptation
In principle, within a system that evolves over time, an ideal general intelligent agent should possess the potential to generate entities for any concept whose fibre can be specified or exemplified. In reality, however, any agent at any single point in time $t$ confronts a finite set of concepts (each with its corresponding fibre), say $K_{t}$. It is this very transition from infinite potential to finite reality that necessitates evaluating intelligence from a dynamic perspective. From this vantage point, intelligence can be assessed along two dimensions: [^2] The terms “synchronic” and “diachronic” are adopted from structural linguistics to distinguish the study of a phenomenon at a particular point in time from its evolution through time, respectively. See Saussure (1916) [22].
@@
    Synchronic Capability: An agent’s breadth and capacity to generate low-$\varepsilon$ entities for the set of concepts $K_{t}$ at a given time $t$.
@@
    Diachronic Capability: The ability of an agent to transition from a state of competence for a set of concepts $K_{t}$ to a subsequent set $K_{t^{\prime}}$. This essentially measures the agent’s adaptive capacity.
Within this framework, concepts such as efficiency, cost, compression, and learning acquire their instrumental value, their importance realized through their potential contribution to these two capabilities. Whether and to what extent an agent leverages these properties stems from the nature of the challenges it confronts—namely, the scale, complexity, and temporal dynamics of the concept sets $K_{t}$.
For instance, when confronted with a large and diverse $K_{t}$, a low-cost generation process might be a strategy for enhancing synchronic capability. Likewise, when concept sets change frequently and unpredictably, a highly compressed knowledge representation could become an advantage for boosting diachronic capability. These are not universal laws, however; the ultimate determinant of a strategy’s merit remains the agent’s overall ability to sustainably generate concept-consistent entities.
6 Generalization to Unseen Concepts
We now consider a training setup based on multiple concepts. Let the training data consist of entities: